Circular No. 7770

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ic IN NGC 7503 W. D. Li, University of California at Berkeley, reports the discovery

by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 17.4) in an unfiltered image taken with the Katzman Automatic Imaging Telescope (KAIT) on Dec. 7.2 UT. The new object is confirmed at about the same brightness on an image taken by M. Schwartz with the Tenagra III 0.5-m telescope on Dec. 8.2 under poor conditions. SN 2001ic is located at α = 23^{h}10^{m}43^{s}.34, δ = +7^{o}34′10′′.4 (equinox 2000.0), which is 14′′.8 east and 6′′.4 north of the nucleus of NGC 7503. A KAIT image taken on Nov. 15.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001id IN UGC 12424 Li also reports the discovery by LOTOSS of another apparent super-

nova (mag ∼ 18.8) in unfiltered KAIT images taken on Dec. 7.2 and 9.2 UT. The new object is located at α = 23^{h}13^{m}10^{s}.88, δ = +10^{o}46′25′′.2 (equinox 2000.0), which is 5′′.7 west and 5′′.6 south of the nucleus of UGC 12424. A KAIT image taken on Nov. 6.2 showed nothing at this position (limiting magnitude ∼ 19.5).

SUPERNOVA 2001hg IN NGC 4162 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001hg (cf. IAUC 7766), obtained by E. Falco on Dec. 7.53 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova before maximum light. The spectrum consists of a very blue continuum with incipient P-Cyg lines of hydrogen and helium.

SUPERNOVA 2001hh IN MCG –02-57-22 Matheson et al. also report that spectra of SN 2001hh (cf. IAUC 7766),

obtained by Falco as above on Dec. 7.04 UT and by Kirshner and Jha on Dec. 7.04 with the Baade 6.5-m telescope at Las Campanas (+ LDSS-2 spectrograph; range 370–1000 nm), show it to be a type-II supernova soon after maximum light. Adopting the NASA/IPAC Extragalactic Database recession velocity of 7445 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 10 000 km/s. Images obtained by Kirshner and Jha on Dec. 6.05 with the Baade 6.5-m telescope yield magnitudes R ∼ 17.8 and I ∼ 17.6.

2001 December 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green